You win $100 in a basketball pool. You have a choice between spending the money now and putting it away for a year in a bank account that pays 5 percent interest. What is the opportunity cost of spending the $100 now


1
Expert's answer
2021-04-12T12:21:36-0400

If the amount of $100 is not spent today and kept in bank account for a year then it will earn a interest of "=\frac{5}{100}\times{100} = 5"

Expected return

"=1000+5=105"

opportunity cost will be the difference between the expected returns of each option.

The opportunity cost of spending the $100 now

"=105-100=5"

$"5"